# Adaptive Depth Networks with Skippable Sub-Paths

### v1

@@ 3.3 Formal Analysis of Skippable Sub-Paths
=$=\mathcal{L}\{\mathbf{h}^{s}_{base}+...+F_{L-1}(\mathbf{h}^{s}_{L-1})+F_{L}(\mathbf{h}^{s}_{L})\}$ (4)
$\approx\mathcal{L}\{\mathbf{h}^{s}_{base}+...+F_{L-1}(\mathbf{h}^{s}_{L-1})\}$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +F_{L}(\mathbf{h}^{s}_{L})\cdot\frac{\partial\mathcal{L}(\mathbf{h}^{s}_{L})}{\partial\mathbf{h}^{s}_{L}}+\mathcal{O}(F_{L}(\mathbf{h}_{L}^{s}))$ (5)
@@
The terms in $\mathcal{O}(F_{L}(\mathbf{h}_{L}^{s}))$ can be ignored if $F_{L}(\mathbf{h}_{L}^{s})$ has a small magnitude. In typical residual networks, however, every layer is trained to learn new features with no constraint, and, hence, there is no guarantee that $F_{L}(\mathbf{h}^{s}_{L})$ have small magnitude. In contrast, in our adaptive depth networks, the residuals in $\mathbf{F}^{s}_{skippable}$ are explicitly enforced to have small magnitude through the skip-aware self-distillation, and, hence, the terms in $\mathcal{O}(F_{L}(\mathbf{h}_{L}^{s}))$ can be ignored for the approximation. If we similarly keep expanding the loss function around $\mathbf{h}_{j}$ ($j=L/2+1,...,L$) while ignoring high order terms, we obtain the following approximation:
@@
Considering this result, we can conjecture that, with our architectural pattern and training method, layers in $\mathbf{F}^{s}_{skippable}$ learn functions that refine input features $\mathbf{h}^{s}_{base}$ iteratively for better inference accuracy while minimally changing the distribution of $\mathbf{h}^{s}_{base}$. Therefore, skipping $\mathbf{F}^{s}_{skippable}$ only slightly reduces prediction accuracy.
3.4 Skip-aware Batch Normalization
@@ 4 Experiments
To demonstrate the generality and effectiveness of our approach for adaptive depth networks, we conduct experiments on various networks and vision tasks.
4.1 Networks
We use four representative residual networks as base models to apply the proposed architecture pattern in Section 3.1: MobileNet V2 [38] is a lightweight CNN model, ResNet [16] is a larger CNN model, and ViT [5] and Swin-T [34] are representative vision transformers. All base models except ViT have 4 residual stages, each with 2 $\sim$ 6 (residual or encoder) blocks. So, according to the proposed architectural pattern, every residual stage is evenly divided into 2 sub-paths for depth adaptation. Since ViT does not define residual stages, we divide 12 encoder blocks into 4 groups, resembling other residual networks, and desiginate the last encoder block of each group as a skippable sub-path.
@@ 4.1 Networks
Table 1: Results on ImageNet. Individual networks with the suffix ‘-Base’ have the same depths as the base-nets of corresponding adaptive depth networks. Latency is measured on an RTX 3090 (batch size = 64).
Model Params FLOPs Acc@1 Latency
(M) (G) (%) (msec)
ResNet50-ADN (super-net) 25.58 4.11 77.6 43.2
ResNet50-ADN (base-net) 2.58 76.1 27.6
ResNet50 (individual) 25.56 4.11 76.7 43.2
ResNet50-Base (individual) 17.11 2.58 75.0 27.6
MV2-ADN (super-net) 3.72 0.32 72.7 15.5
MV2-ADN (base-net) 0.25 70.7 13.2
MV2 (individual) 3.50 0.32 72.1 15.5
MV2-Base (individual) 2.99 0.25 70.7 13.2
ViT-b/16-ADN (super-net) 86.59 17.58 81.3 138.6
ViT-b/16-ADN (base-net) 11.76 79.4 90.7
ViT-b/16 (individual) 86.57 17.58 81.1 138.6
ViT-b/16-Base (individual) 67.70 11.76 78.7 90.7
Swin-T-ADN (super-net) 28.30 4.49 81.8 74.4
Swin-T-ADN (base-net) 2.69 78.7 39.9
Swin-T (individual) 28.29 4.49 81.5 74.4
Swin-T-Base (individual) 15.63 2.69 78.6 39.9
@@ 4.2 ImageNet Classification
We evaluate our method on ILSVRC2012 dataset [37] that has 1000 classes. The dataset consists of 1.28M training and 50K validation images. For CNN models, we follow most training settings in the original papers [16, 38], except that ResNet models are trained for 150 epochs. ViT and Swin-T are trained for 300 epochs, following DeiT’s training recipe [42, 1]. However, in Swin-T-ADN, we disable stochastic depths [23] for the base-net since the strategy of random dropping of residual blocks conflicts with our approach to skipping sub-paths. For fair comparison, our adaptive depth networks and corresponding individual networks are trained in the same training settings. The hyperparameter $\alpha$ in Algorithm 1 is set to $0.5$ for all networks.
@@
The results in Table 1 show that our adaptive depth networks outperform individual counterpart networks even though our sub-networks share parameters in a single network. Further, our results with vision transformers demonstrate that our approach is generally applicable to residual networks and compatible with their state-of-the-art training techniques such as DeiT’s training recipe [42]. We conjecture that this performance improvement results from effective distillation of knowledge from $\textbf{h}^{s}_{super}$ to $\textbf{h}^{s}_{base}$ at each residual stage and the iterative feature refinement at skippable sub-paths, shown in Equation 7.
In Table 2 and Figure 3, several state-of-the-art efficient inference methods and dynamic networks are compared with our sub-networks of adaptive depth networks. The result demonstrates that our adaptive depth networks match or outperform many state-of-the-art static and dynamic networks across varying depth ranges.
@@
In Figure 3, sub-networks of ResNet50-ADN are compared to the counterpart individual ResNets of equivalent depths. In particular, it should be noted that ResNets trained with knowledge distillation in the same training settings has worse performance than individual ResNets trained without knowledge distillation. As reported in previous works, successful knowledge distillation requires a patient and long training [2], and straightforward knowledge distillation using ImageNet does not improve the performance of student models [60, 4]. In contrast, our ResNet50-ADN trained with skip-aware self-distillation achieves better performance than counterpart ResNets. This demonstrates that the high performance of adaptive depth networks does not simply come from the distillation effect, but from the effective combination of the proposed architectural pattern and the skip-aware self-distillation strategy.
Figure 3: Comparison with various approaches on ImageNet. Sub-networks of our adaptive depth networks match or outperform other approaches on ImageNet. ResNet50 (individual) shows individual networks that have equivalent network depths as the sub-networks of ResNet50-ADN. ResNet50 (KD individual) shows individual networks trained using knowledge distillation from pretrained Pytorch ResNet50. In legend, $\dagger$ denotes static compression methods, and $\star$ denotes input-dependent dynamic networks.
@@
Table 2: Our base-nets are compared with state-of-the-art efficient inference methods for various networks. $\dagger$ denotes static compression methods, $\ast$ denotes width-adaptation networks, and $\star$ denotes input-dependent dynamic networks. Note that many of these state-of-the-art methods exploit various non-canonical training techniques, such as iterative retraining and post-processing, while our base-nets are selected from adaptive depth networks without any fine-tuning.
@@
GReg-1 [49$\dagger$ 2.8G 33% 76.1%
DR-ResNet50 ($\alpha$=2.0) [62$\star$ 2.3G 44% 75.3%
DepGraph [7$\dagger$ 2.0G 51% 75.8%
ResNet50-ADN (base-net) 2.6G 37% 76.1%
@@
MV2-ADN (base-net) 0.25G 22% 70.7%
Ada-DeiT-S [35$\star$ 2.3G - 77.3%
WDPruning-0.3-11 [56$\dagger$ 2.6G - 78.3%
DeiT-S+ATS (not fine-tuned) [8$\star$ 3.0G - 78.5%
SPViT-Swin-Ti [28$\star$ 3.3G 27% 80.1%
Swin-T-ADN (base-net) 2.7G 40% 78.7%
@@ 4.3 Training Time
Table 3: Training time (1 epoch), measured on a Nvidia RTX 4090 (batch size = 128). All compared models are based on ResNet50 or have similar FLOPs and depths.
Model Training time (min) # of sub-networks
ResNet50 32.9 1
ResNet50-Base 22.7 1
ResNet50-ADN (ours) 54.7 $2^{4}$ (varying depths)
MSDNet [24] 67.1 9 (early-exit branches)
S-ResNet50 [59] 82.5 4 (varying widths)
SkipNet [50] 90.6 dynamic depths
One important advantage of our approach is that our adaptive depth networks require significantly shorter training time than other adaptive networks. Table 3 shows that training our ResNet50-ADN takes similar time to train two individual networks combined. In contrast, the compared adaptive networks requires much longer training time than ours. MSDNet [24] represents adaptive depth networks with multiple early-exiting branches and classifiers. SkipNet [50] represents input-dependent dynamic networks that drop layers for ‘easy’ inputs. In S-ResNet50 [59], 4 sub-networks with varying widths are trained through iterative self-distillation from its super-net.
Despite significantly shorter training time is required, our adaptive depth networks can support various sub-networks at test time. Figure 4-(left) shows the performance of ResNet50-ADN when its depth is varied at test time. Among them, only the super-net and the base-net are explicitly trained in Algorithm 1. Other sub-networks are selected on the fly at test time by skipping sub-paths in varying number of residual stages. Although these sub-networks are not trained explicitly, they show graceful degradation of performance as the depth of sub-networks becomes gradually shallower. However, although ResNet50-ADN support $2^{4}$ sub-networks in theory, Figure 4-(right) shows that some sub-networks are more useful than the others. For instance, even though ResNet50-ADN(TFFF) and ResNet50-ADN(FFFT) have the same network depths, ResNet50-ADN(TFFF) has about $0.58\%$ higher top-1 accuracy. This result shows that skipping sub-paths in later stages is more detrimental to the performance in general.
Figure 4: Validation accuracy of our sub-networks of varying depths. The boolean values below the bars indicate the residual stages where the sub-paths are skipped.
4.4 Performance on Devices
Figure 5 shows the performance of adaptive networks in an actual device. The inference latency and energy consumption of ResNet50-ADN is compared to S-ResNet50 [59], a representative width-adaptation network. The result shows that depth-adaptation of ResNet50-ADN is highly effective in accelerating inference speeds and reducing energy consumption. For example, in ResNet50-ADN, reducing FLOPs by 38% through depth adaptation reduces both inference latency and energy consumption by 35%. In contrast, even though S-ResNet50 can reduce FLOPs by up to 93% by adjusting its width, it only achieves up to 9% acceleration in practice.
@@ 4.4 Performance on Devices
Figure 5: Inference latency and energy consumption of adaptive networks, measured on Nvidia Jetson Orin Nano (batch size = 1). Although adjusting the network width can support a much wider FLOPs adaptation range, depth adaptation of ResNet50-ADN is more effective for actual inference acceleration and saving energy.
@@
Table 4: Ablation analysis with ResNet50-ADN and ViT-b/32-ADN. Applied components are checked. In the last row, only $D_{KL}$ with softmax outputs is used for the skip-aware self-distillation. $\downarrow$ and $\uparrow$ in parentheses are comparisons to individual networks.
Skip-aware Skip-aware ResNet50 Acc@1 ViT-b/32 Acc@1
self-distillation BNs / LNs Super-net Base-net Super-net Base-net
@@ 4.5 Ablation Study
We conduct ablation study on ImageNet classification to investigate the influence of two key components of the proposed adaptive depth networks: (1) skip-aware self-distillation and (2) skip-aware BN/LNs. When our skip-aware self-distillation is not applied, the loss in Algorithm 1 is modified to $loss=\frac{1}{2}\{criterion(\mathbf{y},\mathbf{\hat{y}}_{super})+criterion(\mathbf{y},\mathbf{\hat{y}}_{base})\}$ for joint training of the super-net and the base-net. Table 4 shows the results. For ResNet50, when neither of them is applied, the inference accuracy of the super-net and the base-net is significantly lower than individual networks by 1.5% and 2.8%, respectively. This result shows the difficulty of joint training sub-networks for adaptive networks. When one of the two components is applied individually, the performance is still slightly worse than individual networks’. In the third row, when both skip-aware self-distillation and skip-aware BNs are applied together, ResNet50-ADN achieves significantly better performance than individual networks, both in the super-net and the base-net. Finally, the last row shows the result when only $D_{KL}(\mathbf{\hat{y}}_{super}\|\mathbf{\hat{y}}_{base})$ is used for skip-aware self-distillation. Without explicit distillation of intermediate features, slightly lower performance is observed in the base-net. The results with ViT-b/32 show that switchable layer normalization has similar effect in vision transformers.
(a) Original ResNet50
(b) ResNet50-ADN (super-net)
(c) ResNet50-ADN (base-net)
Figure 6: Class Activation Maps of the 3rd residual stages of ResNet50s. (a) Original ResNet50’s activation regions change gradually across all blocks. (b) In ResNet50-ADN (super-net), the first 3 blocks have extensive hot activation regions, implying active learning of new level features. In contrast, the skippable last 3 blocks have far less activation regions and they are gradually refined around the target. (c) The base-net’s final activation map is very similar to super-net’s final activation map, implying that they are at the same feature level.
4.6 Visual Analysis of Sub-Paths
To investigate how our training method affects feature representations in the mandatory and the skippable sub-paths, we visualize the activation of 3rd residual stage of ResNet50-ADN using Grad-CAM [39]. The 3rd residual stage of ResNet50-ADN has 6 residual blocks and the last three blocks are skippable. In Figure 6-(a), the activation regions of original ResNet50 changes gradually across all consecutive blocks. In contrast, in Figure 6-(b), ResNet50-ADN (super-net) manifests very different activation regions in two sub-paths. In the first three residual blocks, we can observe lots of hot activation regions in wide areas, suggesting active learning of new level features. In contrast, significantly less activation regions are found in the skippable last three blocks and they are gradually concentrated around the target object, demonstrating the refinement of learned features. Further, in Figure 6-(c), we can observe that the final activation map of the ResNet50 (base-net) is very similar to the super-net’s final activation map in Figure 6-(b). This implies that they have similar distributions for the same inputs, as suggested in Section 3.3.
4.7 Object Detection and Instance Segmentation

### v2

@@ 3.3 Analysis of Skippable Sub-Paths
=$=\mathcal{L}\{\mathbf{h}^{s}_{base}+F_{L/2+1}(\mathbf{h}^{s}_{L/2+1})+...+F_{L-1}(\mathbf{h}^{s}_{L-1})+F_{L}(\mathbf{h}^{s}_{L})\}$ (3)
$\approx\mathcal{L}\{\mathbf{h}^{s}_{base}+F_{L/2+1}(\mathbf{h}^{s}_{L/2+1})+...+F_{L-1}(\mathbf{h}^{s}_{L-1})\}+F_{L}(\mathbf{h}^{s}_{L})\cdot\frac{\partial\mathcal{L}(\mathbf{h}^{s}_{L})}{\partial\mathbf{h}^{s}_{L}}+\mathcal{O}(F_{L}(\mathbf{h}_{L}^{s}))$ (4)
@@
The high-order terms in $\mathcal{O}(F_{L}(\mathbf{h}_{L}^{s}))$ can be ignored if $F_{L}(\mathbf{h}_{L}^{s})$ has a small magnitude. In typical residual networks, however, every layer is trained to learn new features with no constraint, and, hence, there is no guarantee that $F_{L}(\mathbf{h}^{s}_{L})$ have small magnitude. In contrast, in our adaptive depth networks, the residuals in $\mathbf{F}^{s}_{skippable}$ are explicitly enforced to have small magnitude through the proposed self-distillation strategy, and, hence, the terms in $\mathcal{O}(F_{L}(\mathbf{h}_{L}^{s}))$ can be ignored for the approximation; see Figure 3 for empirical evidence. If we similarly keep expanding the loss function around $\mathbf{h}^{s}_{j}$ ($j=L/2+1,...,L$) while ignoring high order terms, we obtain the following approximation:
@@
Considering this result, we can conjecture that, with our architectural pattern and self-distillation strategy, layers in $\mathbf{F}^{s}_{skippable}$ learn functions that refine input features $\mathbf{h}^{s}_{base}$ iteratively for better inference accuracy while minimally changing the distribution of $\mathbf{h}^{s}_{base}$.
Figure 3: $||F(\mathbf{h})||_{2}/||\mathbf{h}||_{2}$ at residual blocks. In ours, skippable sub-paths (orange areas) minimally change the distribution of input h
Empirical Analysis: We can estimate how much the distribution of input $\mathbf{h}$ is transformed by residual function $F$ by measuring $||F(\mathbf{h})||_{2}/||\mathbf{h}||_{2}$ at each residual block. Figure 3 illustrates $||F(\mathbf{h})||_{2}/||\mathbf{h}||_{2}$ at every residual block in the baseline ResNet50 and our ResNet50-ADN that is trained according to Algorithm 1  [^2] ImageNet validation dataset is used for this experiment.. Our ResNet50-ADN exhibits greater transformation than ResNet50 in the mandatory sub-paths (blue areas) and less transformation in the skippable sub-paths (orange areas). This result demonstrates that our self-distillation strategy in Algorithm 1 effectively trains the skippable sub-paths to minimally change the input distribution. As a result, the blocks in skippable sub-paths can be skipped with less impact on performance.
3.4 Skip-Aware Batch/Layer Normalization
@@ 4 Experiments
We use four representative residual networks both from CNNs and vision transformers as base models to apply the proposed architecture pattern: MobileNet V2 [35] is a lightweight CNN model, ResNet [34] is a larger CNN model, and ViT [37] and Swin-T [38] are representative vision transformers. All base models except ViT have hierarchical stages, each with 2 $\sim$ 6 residual blocks. So, according to the proposed architectural pattern, every residual stage is evenly divided into 2 sub-paths for depth adaptation. ViT does not define hierarchical stages and all residual encoder blocks have same spatial dimensions. Therefore, we divide 12 encoder blocks into 4 groups, resembling other residual networks, and select the last encoder block of each group as a skippable sub-path. Details are in Table 4 in Appendix A.1.
@@
Model Params FLOPs Acc
(M) (G) (%)
ResNet50-ADN (FFFF) 25.58 4.11 77.6
ResNet50-ADN (TTTT) 2.58 76.1
ResNet50 25.56 4.11 76.7
ResNet50-Base 17.11 2.58 75.0
MbV2-ADN (FFFF) 3.72 0.32 72.7
MbV2-ADN (TTTT) 0.25 70.7
MbV2 3.50 0.32 72.1
MbV2-Base 2.99 0.25 70.7
ViT-b/16-ADN (FFFF) 86.59 17.58 81.3
ViT-b/16-ADN (TTTT) 11.76 79.4
ViT-b/16 86.57 17.58 81.1
ViT-b/16-Base 67.70 11.76 78.7
Swin-T-ADN (FFFF) 28.30 4.49 81.8
Swin-T-ADN (TTTT) 2.69 78.7
Swin-T 28.29 4.49 81.5
Swin-T-Base 15.63 2.69 78.6
Figure 4: (a) Results on ImageNet validation dataset. Networks with the suffix ‘-Base’ have the same depths as the base-nets of corresponding adaptive depth networks. (b) Pareto frontiers formed by the sub-networks of our adaptive depth networks. ResNet50 (individual) and ResNet50 (KD individual) are non-adaptive networks having same depths as the sub-networks of ResNet50-ADN.
@@ 4.1 ImageNet Classification
We evaluate our method on ILSVRC2012 dataset [49] that has 1000 classes. The dataset consists of 1.28M training and 50K validation images. For CNN models, we follow most training settings in the original papers [34, 35], except that ResNet models are trained for 150 epochs. ViT and Swin-T are trained for 300 epochs, following DeiT’s training recipe [50, 51]. For Swin-T-ADN, we disable stochastic depths [40] for the mandatory sub-paths since the strategy of random dropping of residual blocks conflicts with our approach to skipping sub-paths. For fair comparison, our adaptive depth networks and corresponding individual networks are trained in the same training settings.
@@
The results in Figure 4-(a) show that our adaptive depth networks outperform counterpart individual networks even though many sub-networks share parameters in a single model. Further, our results with vision transformers demonstrate that our approach is generally applicable and compatible with their state-of-the-art training techniques such as DeiT’s training recipe [50]. We conjecture that this performance improvement results from effective distillation of knowledge from $\textbf{h}^{s}_{super}$ to $\textbf{h}^{s}_{base}$ at each residual stage and the iterative feature refinement at skippable sub-paths, shown in Equation 6.
@@
Figure 4-(b) shows Pareto frontiers formed by selected sub-networks of our adaptive depth networks; Table 5 in Appendix A.2 shows the performance of all sub-networks. In Figure 4-(b) and Table 1, several state-of-the-art efficient inference methods and dynamic networks are compared with our base-networks. The result demonstrates that our adaptive depth networks show competitive performance across a range of FLOPs. In Figure 4-(b), it should be noted that individual ResNets trained with knowledge distillation has worse performance than individual ResNets. As reported in previous works, successful knowledge distillation requires a patient and long training [52], and straightforward knowledge distillation using ImageNet does not improve the performance of student models [53, 54]. In contrast, our ResNet50-ADN trained with the proposed self-distillation strategy achieves better performance than counterpart ResNets. This demonstrates that the high performance of adaptive depth networks does not simply come from distillation effect.
@@
Table 1: Our base-nets are compared with state-of-the-art efficient inference methods. $\dagger$ denotes static pruning methods, $\ast$ denotes width-adaptation networks, and $\star$ denotes input-dependent dynamic networks. While these approaches exploit various non-canonical training techniques, such as iterative retraining, our base-nets are instantly selected from adaptive depth networks without fine-tuning.
@@
GReg-1 [55$\dagger$ 2.8G 33% 76.1%
DepGraph [56$\dagger$ 2.0G 51% 75.8%
DR-ResNet50 ($\alpha$=2.0) [47$\star$ 2.3G 44% 75.3%
ResNet50-ADN (TTTT) 2.6G 37% 76.1%
@@
MbV2-ADN (TTTT) 0.25G 22% 70.7%
WDPruning-0.3-11 [57$\dagger$ 2.6G - 78.3%
X-Pruner [58] $\dagger$ 3.2G 28% 80.7%
Ada-DeiT-S [27$\star$ 2.3G - 77.3%
SPViT-Swin-Ti [59$\star$ 3.3G 27% 80.1%
Swin-T-ADN (TTTT) 2.7G 40% 78.7%
@@ 4.2 Training Time
One major advantage of our adaptive depth networks is that they require significantly shorter training time than prior adaptive networks. Figure 5-(a) shows that training our adaptive depth networks take similar time to train two individual networks combined. This is because our training method in Algorithm 1 exclusively exploits only the super-net and the base-net for self-distillation. In contrast, compared adaptive networks require much longer training time than ours since they have to train every target sub-network explicitly  [^3] Although the prior works we compared for training time may seem outdated, they remain relevant and representative because there has been little progress in reducing the training time of adaptive networks.. For example, on every training iteration, AlphaNet [11] randomly samples sub-networks from its search space for self-distillation. While AlphaNet [11] supports significantly more sub-networks, it should be noted that supporting many sub-networks is not our goal. Instead, our objective is to provide better performance Pareto with several useful sub-networks, as demonstrated in Figure 4-(b). For example, although ResNet50-ADN can select $2^{4}$ sub-networks, Table 5 (in Appendix A.2) shows that some sub-networks are more useful than the others; results in Table 5 show that skipping sub-paths in deeper residual stages is more detrimental to the performance in general, implying that refinement of features is more needed for deeper residual stages.
Model 1 epoch (min) # of sub-nets
ResNet50 32.9 -
ResNet50-Base 22.7 -
ResNet50-ADN (ours) 54.7 $2^{4}$ depths
MSDNet [8] 67.1 9 depths
S-ResNet50 [10] 82.5 4 widths
MbV2 13.1 -
MbV2-Base 10.0 -
MbV2-ADN (ours) 22.5 $2^{4}$ depths
AlphaNet [11] 230.5 216 depths
@@
Figure 5: (a) Training time (1 epoch), measured on Nvidia RTX 4090 (batch size: 128). AlphaNet is configured to have similar FLOPs to MbV2 and makes sub-networks by only adjusting the network depth. (b) Inference latency and energy consumption of adaptive networks, measured on Nvidia Jetson Orin Nano (batch size: 1)
4.3 On-Device Performance
While the Pareto frontiers formed by sub-networks demonstrate theoretic performance, inference acceleration in actual devices is more important in practice for effective control of inference latency and energy consumption.
Figure 5(b) shows the performance in Nvidia Jetson Orin Nano. The inference latency and energy consumption of ResNet50-ADN is compared to S-ResNet50, a representative width-adaptation network. The result shows that depth-adaptation of ResNet50-ADN is highly effective in accelerating inference speeds and reducing energy consumption. For example, in ResNet50-ADN, reducing FLOPs by 38% through depth adaptation reduces both inference latency and energy consumption by 35%. In contrast, even though S-ResNet50 can reduce FLOPs by up to 93% by adjusting its width, it only achieves up to 9% acceleration in practice.
4.4 Ablation Study
@@
Table 2: Ablation analysis with ResNet50-ADN and ViT-b/32-ADN. Applied components are checked. $\downarrow$ and $\uparrow$ in parentheses are comparisons to non-adaptive individual networks.
self- skip-aware ResNet50 Acc@1 (%) ViT-b/32 Acc@1 (%)
distllation. BNs/LNs FFFF TTTT FFFF TTTT
@@
We first investigate the influence of two key components of the proposed adaptive depth networks: (1) self-distillation of sub-paths and (2) skip-aware BNs/LNs. When our self-distillation method is not applied, the loss of the base-net, or $loss_{base}$, in Algorithm 1 is modified to $criterion(\mathbf{y},\mathbf{\hat{y}}_{base})$. Table 2 shows the results. For ResNet50, when neither of them is applied, the inference accuracy of the super-net and the base-net is significantly lower than non-adaptive individual networks by 1.5% and 2.8%, respectively. This result shows the difficulty of joint training sub-networks for adaptive networks. When one of the two components is applied individually, the performance is still slightly worse than individual networks’. In the last row, when both self-distillation and skip-aware BNs are applied together, ResNet50-ADN achieves significantly better performance than individual networks, both in the super-net and the base-net. The results with ViT-b/32 show that switchable layer normalization has similar effect in vision transformers.
Table 3: Comparison of self-distillation strategies. Our approach (in bold) uses exclusively the super-net and the base-net as a teacher and a student, respectively.
sub-nets ResNet50-ADN Acc@1 (%) ViT/b-32-ADN Acc@1 (%)
Teacher Student FFFF TFFF TTFF TTTF TTTT FFFF TFFF TTFF TTTF TTTT
FFFF TTTT 77.6 77.2 76.8 76.1 76.1 76.6 76.0 75.5 74.5 74.3
FFFF Random 77.1 76.7 76.4 75.5 74.8 75.2 74.7 73.9 72.9 71.1
Random TTTT 75.5 75.5 75.4 75.0 74.9 72.0 72.0 71.9 71.8 71.7
Random Random 75.4 75.2 75.2 74.9 74.6 70.8 70.7 70.6 70.4 70.3
Self-Distillation Strategies: Our self-distillation approach in Algorithm 1 exploits only two sub-networks exclusively as a teacher and a student. Specifically, the super-nets, or FFFF, acts as the the teacher and the base-nets, or TTTT, becomes the student. The purpose of this strategy is not to train only those two sub-networks, but rather to train skippable sub-paths in a way that minimally modifies the feature distribution, as demonstrated in Figure 3. To investigate the effect of self-distillation strategies, we conduct an experiment in Table 3. In every training iteration, rather than exclusively using FFFF and TTTT sub-networks for self-distillation, we randomly sample either a teacher, a student, or both from $2^{4}$ sub-networks. These randomly sampled sub-networks are trained explicitly through self-distillation. However, as shown in Table 3, all sub-networks, both from ResNet50-ADN and ViT/b-32-ADN, perform significantly better when our self-distillation strategy is applied for training. Even though some sub-networks (such as TTTF, TTFF, and TTTF) are not trained explicitly with our self-distillation strategy, they still outperform their counterpart sub-networks trained through random sampling. This result demonstrates that our method of training sub-paths is more effective than training target sub-networks themselves.
Figure 6: Pareto frontiers of ResNet50-ADNs trained with varying ratios between mandatory and skippable sub-paths. Total number of blocks remains unchanged.
Lengths of Mandatory Sub-Paths: In Figure 6, we investigate the effect of varying the length ratio between the mandatory and the skippable sub-paths. (Details are in Table 6.) If mandatory sub-paths become shorter than skippable sub-paths, smaller sub-networks can be selected since more layers can be skipped. However, the result in Figure 6 shows that this configuration significantly degrades performance of all sub-networks. Since every sub-network shares parameters of mandatory sub-paths, low inference capability of shallow mandatory sub-paths affects all sub-networks. This implies that maintaining certain depths in mandatory sub-paths is essential for effective inference. In contrast, increasing the length of mandatory sub-paths further does not yield further performance improvement and only reduce the range of FLOPs adaptation.
5 Conclusions
We propose a practical approach to adaptive depth networks that can be applied to both CNNs and transformers with minimal training effort. We provide a general principle and formal explanation on how depth adaptation can be achieved with minimal performance degradation. Under this principle, our approach can avoid typical exhaustive training of target sub-network and instead focus on optimizing the sub-paths of the network to have specific properties. At test time, those sub-paths can be connected in a combinatorial manner to construct sub-networks with various accuracy-efficiency trade-offs from a single network. Experimental results show that those sub-networks form a better Pareto frontier than non-adaptive baseline networks and achieve actual inference acceleration.
Appendix A Appendix: Detailed Settings and Evaluation Results
A.1 Detailed Architectures
number of number of
mandatory skippable total
blocks blocks blocks
stage 1 2 1 3
stage 2 2 2 4
stage 3 3 3 6
stage 4 2 1 3
(a) ResNet50-ADN
number of number of
mandatory skippable total
blocks blocks blocks
stage 1 2 1 3
stage 2 2 1 3
stage 3 2 1 3
stage 4 2 1 3
(b) Vit-b/16-ADN
number of number of
mandatory skippable total
blocks blocks blocks
stage 1 1 - 1
stage 2 2 - 2
stage 3 2 1 3
stage 4 2 2 4
stage 5 2 1 3
stage 6 2 1 3
stage 7 1 - 1
(c) MbV2-ADN
number of number of
mandatory skippable total
blocks blocks blocks
stage 1 1 1 2
stage 2 1 1 2
stage 3 3 3 6
stage 4 2 - 2
(d) Swin-T-ADN
Table 4: Each residual stage of base models is evenly divided into two sub-paths; the first is mandatory and the other is skippable. Since ViT does not define hierarchical stages, 12 identical encoder blocks are divided into 4 stages, resembling other residual networks for vision tasks.
A.2 Performance of Sub-Networks
sub-network FLOPs (G) Acc@1 (%)
FFFF 4.11 77.6
TFFF 3.90 77.2
FTFF 3.68 76.5
FFTF 3.46 75.6
FFFT 3.90 76.2
TTFF 3.46 76.8
TFTF 3.24 75.4
TFFT 3.68 76.1
FTTF 3.02 75.3
FTFT 3.46 75.4
FFTT 3.25 75.6
TTTF 2.80 76.1
TTFT 3.24 75.9
TFTT 3.02 75.4
FTTT 2.80 75.6
TTTT 2.58 76.1
(a) ResNet50-ADN
sub-network FLOPs (G) Acc@1 (%)
FFFF 17.58 81.3
TFFF 16.20 80.5
FTFF 16.20 80.2
FFTF 16.20 80.0
FFFT 16.20 80.4
TTFF 14.67 79.9
TFTF 14.67 79.6
TFFT 14.67 79.8
FTTF 14.67 79.4
FTFT 14.67 79.7
FFTT 14.67 79.4
TTTF 13.21 79.3
TTFT 13.21 79.4
TFTT 13.21 79.2
FTTT 13.21 79.1
TTTT 11.76 79.1
(b) ViT-b/16-ADN
Table 5: FLOPs and ImageNet validation accuracy of sub-networks. Only the super-net and the base-net are trained explicitly. The other parameter-sharing sub-networks in the middle can be selected at test time without explicit training. The highest accuracy in each group is shown in bold. This results show that skipping sub-paths in deeper residual stages is more detrimental to the performance in general, implying that refinement of features is more needed for deeper residual stages.
A.3 Varying the Lengths of Sub-Paths in ResNet50-ADN
number of number of total
mandatory skippable blocks
blocks blocks
stage 1 1 2 3
stage 2 1 3 4
stage 3 2 4 6
stage 4 1 2 3
(a) # of mandatory $<$ # of skippable
number of number of total
mandatory skippable blocks
blocks blocks
stage 1 2 1 3
stage 2 3 1 4
stage 3 4 2 6
stage 4 2 1 3
(b) # of mandatory $>$ # of skippable
Table 6: Varying the length ratio of sub-paths of ResNet50-ADN. Total number of blocks of each stage is not changed.
A.4 Varying Sub-Paths Length Ratios in Vit-b/16-ADN
Each stage of Vit-b/16-ADN has 3 encoder blocks and, by default, we select every last encoder block as a skippable sub-path. Therefore every stage has two mandatory blocks and 1 skipable blocks as shown in Table 4-(b). Figure 7-(a) shows different configuration where every last two blocks of the stages become skippable. With this configuration in Figure 7-(a), we can select much smaller sub-networks. For example, the smallest sub-network, or the base-net, of Vit-b/15-ADN has only 4 mandatory blocks and it requires 5.82 GFLOPs. However, the result in Figure 7-(b) shows that this configuration significantly degrades performance of all sub-networks. As demonstrated with ResNet50-ADN in Figure 6, low inference capability of shallow mandatory sub-paths affects all sub-networks. This again shows that some minimum depths in mandatory sub-paths is required for effective inference.
number of number of total
mandatory skippable blocks
blocks blocks
stage 1 1 2 3
stage 2 1 2 3
stage 3 1 2 3
stage 4 1 2 3
Figure 7: (a) The configuration of Vit-b/16-ADN with longer skippable sub-paths. (b) Pareto-frontier when different length ratios between the mandatory and the skippable sub-paths are applied.
Appendix B Appendix: More Experiments and Analysis
B.1 Object Detection and Instance Segmentation